OpenStudy (anonymous):

sqrt(1/((3/x)+4) +4) +3)=?

OpenStudy (anonymous):

\[\sqrt{\frac{1}{\frac{3}{x}+4+4+3}}\]Or this?

OpenStudy (anonymous):

\[\sqrt{\frac{1}{\frac{3}{x}+11}}\]\[\sqrt{\frac{1}{\frac{3}{x}+\frac{11}{x}}}\]Continue as far as you can, call me for help if needed

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Got the answer?